OpenStudy (anonymous):

Solve the equation 4x=8/2 = −10 for x

jimthompson5910 (jim_thompson5910):

You have too many equal signs

OpenStudy (anonymous):

oh its addition

jimthompson5910 (jim_thompson5910):

So 4x+8/2 = −10 ??

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

So it really looks like this \[\Large \frac{4x+8}{2} = -10\] OR this \[\Large 4x+\frac{8}{2} = -10\] ???

OpenStudy (anonymous):

the first one.

jimthompson5910 (jim_thompson5910):

\[\Large \frac{4x+8}{2} = -10\] \[\Large 4x+8 = -10*2\] \[\Large 4x+8 = -20\] Can you take it from here? If so, tell me what you get.

OpenStudy (anonymous):

yes but what happens to the 2?

jimthompson5910 (jim_thompson5910):

I multiplied both sides by 2 (to get rid of it on the left side)

OpenStudy (anonymous):

oh okay.